OpenStudy (okdutchman7):

Please help. WILL MEDAL! Adrian went to the fruit store because they were selling p pears for c cents. When he arrived, the store announced a special of 5 cents less on each pear. How many pears could Adrian buy for 4 dollars? @jim_thompson5910 @Mehek14 @nincompoop @sleepyjess @satellite73 @zepdrix @Loser66 @just_one_last_goodbye @Elsa213

OpenStudy (okdutchman7):

\[\frac{ c }{ p }\frac{ cents }{ pear }\]

OpenStudy (okdutchman7):

^original price

OpenStudy (okdutchman7):

\[\frac{ c-5 }{ p } \frac{ cents }{ pear }\]

OpenStudy (okdutchman7):

^new price

OpenStudy (okdutchman7):

This is where I get stuck.

OpenStudy (okdutchman7):

It is an abstract equation so you won't find an "real" number.

OpenStudy (okdutchman7):

It will contain algebraic variables.

jimthompson5910 (jim_thompson5910):

@okdutchman7 the cost of one pear before the discount is c/p cents after you apply the 5 cent discount for each pear, then you end up with the cost (c/p) - 5 cents for each pear

OpenStudy (okdutchman7):

good catch jim i read that wrong

OpenStudy (okdutchman7):

\[\frac{ c }{ p }-5\frac{ cents }{ pear }\]

jimthompson5910 (jim_thompson5910):

Example: c = 100 p = 2 c/p = 100/2 = 50 cents for one pear (c/p)-5 = 50-5 = 45 cents per pear

OpenStudy (okdutchman7):

I understand that.

OpenStudy (okdutchman7):

I don't know how to relate dollars and cents.

OpenStudy (okdutchman7):

since I don't know how many cents c represents

jimthompson5910 (jim_thompson5910):

Let x = number of pears bought y = cost per pear (after discount) z = total cost All costs are in cents In this case, x = unknown y = (c/p) - 5 cents per pear z = 400 cents That means x*y = z x*((c/p) - 5) = 400 solve for x

OpenStudy (okdutchman7):

\[x=\frac{ 400 }{ c-5p }\]

OpenStudy (okdutchman7):

What does this^ mean?

jimthompson5910 (jim_thompson5910):

\[\Large x*\left(\frac{c}{p}-5\right) = 400\] \[\Large \frac{x*\left(\frac{c}{p}-5\right)}{\frac{c}{p}-5} = \frac{400}{\frac{c}{p}-5}\] \[\Large x = \frac{400}{\frac{c}{p}-5}\] \[\Large x = \frac{400}{\frac{c}{p}-5}\times\frac{p}{p}\] \[\Large x = \frac{400p}{\left(\frac{c}{p}-5\right)*p}\] \[\Large x = \frac{400p}{\left(\frac{c}{p}\right)*p-\left(5\right)*p}\] \[\Large x = \frac{400p}{c-5p}\] So you are close but not quite there

OpenStudy (okdutchman7):

That's what I got the first time :)

jimthompson5910 (jim_thompson5910):

What does it mean? It represents the number of pears you can buy for \(\Large \frac{c}{p}-5\) cents if you only have 400 cents (4 dollars)

jimthompson5910 (jim_thompson5910):

you forgot the p with the 400 though

OpenStudy (okdutchman7):

Yeah that's what I got the first time.

OpenStudy (okdutchman7):

\[\frac{ 400p }{ c-5p }\]

OpenStudy (okdutchman7):

Is that the final answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (okdutchman7):

Thanks!

jimthompson5910 (jim_thompson5910):

np